OpenStudy (anonymous):

2 questions. 1) -1<= a - 3 <= 2 2)A towns high temperature for a month is 88F and the lowest temperature is 65F. Write a compound inequality to represent the range of temperatures for the given month.

OpenStudy (anonymous):

\[-1+3\le a \le 2+3\] \[2 \le a \le 5\] 2)\[65T \le T \le 88F\]